The sum of two counting numbers is 91. Half of one number equals three times the other number. What are the number?
Answer by checkley71(8403)

X+Y=91 OR X=91-Y
X/2=3Y
(91-Y)/2=3Y NOW CROSS MULTIPLY
6Y=91-Y
7Y=91
Y=91/7
Y=13 ANSWER.
X+13=91
X=91-13
X=78 ANSWER.
PROOF:
78/2=3*13
39=39